Rectangular building: A rectangular building has a perimeter of 1200 feet. (a) If one side of the building has length L, write a polynomial expression that gives its area. (be sure to multiply your expression.)(b) Evaluate the expression in part (a) for L=50 and interpret the answer.
Answer by stanbon(75887)

Rectangular building: A rectangular building has a perimeter of 1200 feet.
(a) If one side of the building has length L, write a polynomial expression that gives its area. (be sure to multiply your expression.)
-----
P = 2L+2W
600 = L+W
----
Area = L(600-L)
Area = 600L-L^2
==========================
(b) Evaluate the expression in part (a) for L=50 and interpret the answer.
Area = 600*50-50^2
Area = 3000-2500
Area = 500 sq.ft.
====================
Cheers,
Stan H.